\begin{thebibliography}{10}

\bibitem{E:ZZS:73:Praparata}
G.~V. Zaitsev, V.~A. Zinoviev, and N.~V. Semakov.
\newblock Interrelation of {P}reparata and {H}amming codes and extension of
  {H}amming codes to new double-error-correcting codes.
\newblock In P.~N. Petrov and F.~Csaki, editors, {\em Proc. 2nd Int. Symp.
  Information Theory, Tsahkadsor, Armenia, USSR, 1971}, pages 257--264,
  Budapest, Hungary, 1973. Akademiai Kiado.

\bibitem{E:Preparata:1968}
F.~P. Preparata.
\newblock A class of optimum nonlinear double-error correcting codes.
\newblock {\em
  \href{http://www.sciencedirect.com/science/journal/00199958}{Inf. Control}},
  13(4):378--400, 1968.
\newblock \DOI{10.1016/S0019-9958(68)90874-7}.

\bibitem{E:Vas:nongroup_perfect}
Yu.~L. Vasil'ev.
\newblock On nongroup close-packed codes.
\newblock In {\em Problemy Kibernetiki}, volume~8, pages 337--339. 1962.
\newblock In Russian, English translation in Probleme der Kybernetik, 8: 92-95,
  1965.

\bibitem{E:Sol:2008:survey}
F.~I. Solov'eva.
\newblock On perfect binary codes.
\newblock {\em
  \href{http://www.sciencedirect.com/science/journal/0166218X}{Discrete Appl.
  Math.}}, 156(9):1488--1498, 2008.
\newblock \DOI{10.1016/j.dam.2005.10.023}.

\bibitem{E:Hed:2008:survey}
O.~Heden.
\newblock A survey of perfect codes.
\newblock {\em \href{http://aimsciences.org/journals/amc/}{Adv. Math.
  Commun.}}, 2(2):223--247, 2008.
\newblock \DOI{10.3934/amc.2008.2.223}.

\bibitem{E:Lloyd}
S.~P. Lloyd.
\newblock Binary block coding.
\newblock {\em \href{http://www.alcatel-lucent.com/bstj/}{Bell Syst. Tech.
  J.}}, 36(2):517--535, 1957.
\newblock \DOI{10.1002/j.1538-7305.1957.tb02410.x}.

\bibitem{E:ShSl}
H.~S. Shapiro and D.~L. Slotnick.
\newblock On the mathematical theory of error correcting codes.
\newblock {\em
  \href{http://ieeexplore.ieee.org/xpl/RecentIssue.jsp?punumber=5288520}{IBM J.
  Res. Dev.}}, 3(1):25--34, 1959.

\bibitem{E:Dumer:1976}
I.~I. Dumer.
\newblock Some new uniformly packed codes.
\newblock Proc. Moscow Inst. Physics and Technology. Ser. ``Radiotekhnika i
  Elektronika'', pages 72--78. 1976.
\newblock in Russian.  
  
\bibitem{E:BvLW}
R.~D. Baker, J.~H. van Lint, and R.~M. Wilson.
\newblock On the {P}reparata and {G}oethals codes.
\newblock {\em
  \href{http://ieeexplore.ieee.org/xpl/RecentIssue.jsp?punumber=18}{IEEE Trans.
  Inf. Theory}}, 29(3):342--345, 1983.
\newblock \DOI{10.1109/TIT.1983.1056675}.

\bibitem{E:vDamFDFla:2000}
E.~R. van Dam and D.~Fon-Der-Flaass.
\newblock Uniformly packed codes and more distance regular graphs from crooked
  functions.
\newblock {\em \href{http://link.springer.com/journal/10801}{J. Algebr.
  Comb.}}, 12(2):115--121, 2000.
\newblock \DOI{10.1023/A:1026583725202}.

\bibitem{E:HammonsOth:Z4_linearity}
A.~R. Hammons, Jr, P.~V. Kumar, A.~R. Calderbank, N.~J.~A. Sloane, and
  P.~Sol\'e.
\newblock The {$Z_4$}-linearity of {K}erdock, {P}reparata, {G}oethals, and
  related codes.
\newblock {\em
  \href{http://ieeexplore.ieee.org/xpl/RecentIssue.jsp?punumber=18}{IEEE Trans.
  Inf. Theory}}, 40(2):301--319, 1994.
\newblock \DOI{10.1109/18.312154}.

\bibitem{E:Kantor:2004symplecticsemifield}
W.~M. Kantor and M.~E. Williams.
\newblock Symplectic semifield planes and {$Z_4$}-linear codes.
\newblock {\em \href{http://www.ams.org/tran/}{Trans. Am. Math. Soc.}},
  356(3):895--938, 2004.
\newblock \DOI{10.1090/S0002-9947-03-03401-9}.

\bibitem{E:MWS}
F.~J. MacWilliams and N.~J.~A. Sloane.
\newblock {\em The Theory of Error-Correcting Codes}.
\newblock Amsterdam, Netherlands: North Holland, 1977.

\bibitem{SZZ:1971:UPC}
N.~V. Semakov, V.~A. Zinoviev, and G.~V. Zaitsev.
\newblock Uniformly packed codes.
\newblock {\em \href{http://link.springer.com/journal/11122}{Probl. Inf.
  Transm.}}, 7(1):30--39, 1971.
\newblock Translated from \em Probl. Peredachi Inf.\em, 7(1): 38-50, 1971.

\bibitem{E:Sol88en}
F.~I. Solov'eva.
\newblock On the factorization of code-generating d.n.f.
\newblock In {\em Metody diskretnogo analiza v issledovanii funktsionalnyh
  sistem}, volume~47, pages 66--88. Institute of Mathematics of SB AS USSR,
  Novosibirsk, 1988.
\newblock In Russian.

\bibitem{E:Tok:2004.en}
N.~N. Tokareva.
\newblock On components of {P}reparata codes.
\newblock {\em \href{http://link.springer.com/journal/11122}{Probl. Inf.
  Transm.}}, 40(2):159--164, 2004.
\newblock \DOI{10.1023/B:PRIT.0000043933.08204.89}, translated from
  \href{http://www.mathnet.ru/php/journal.phtml?jrnid=ppi\&option_lang=eng}{Probl.
  Peredachi Inf.} 40(2) 2004, 63--69.

\end{thebibliography}

\documentclass[11pt,russian]{article}   
\usepackage{cmap}
\usepackage[T2A]{fontenc}
\usepackage[utf8]{inputenc} 
\usepackage[english]{babel}
\usepackage{amsthm}
\normalfont

\usepackage{amssymb,amsmath}
\usepackage{color,graphicx}

\makeatletter
\def\@seccntformat#1{\csname the#1\endcsname.\ } 
\makeatother

\newtheorem{theorem}{Theorem}
\newtheorem{conj}{Conjecture}

\begin{document}

\noindent
\textbf{\LARGE On perfect codes that do not include \\ Preparata-like codes%
\footnote{The work was funded by the Russian science foundation (grant 14-11-00555)}}
\date{}

\vspace*{5mm} \noindent
\textsc{D. S. Krotov, A. Yu. Vasil'eva} \hfill \texttt{krotov,vasilan@math.nsc.ru} \\
{\small Sobolev Institute of Mathematics, Novosibirsk, Russia} \\

\medskip

\begin{center}
\parbox{11,8cm}{\footnotesize
\textbf{Abstract.}
We show that for every length of form $4^k-1$,
there exists a binary $1$-perfect code that does not include any Preparata-like code.}
\end{center}

\baselineskip=0.9\normalbaselineskip

\section{Introduction}

The binary $1$-error-correcting perfect codes (below, perfect codes)
and the codes with the parameters of the Preparata codes
(below, Preparata codes)
are two remarkable infinite families of binary codes having many common properties.
Many of these properties are related to the fact that
any code with the parameters (length, size, code distance)
of the codes from the considered classes
induces a regular partition of the space.
Every Preparata code is included in some perfect code \cite{E:ZZS:73:Praparata}.

A perfect code can include several nonequivalent Preparata codes.
In the current correspondence, we consider the question
of the existence of perfect codes that does not include any Preparata code
(of course, we consider only lengths of form $4^n-1$,
for which Preparata codes exist \cite{E:Preparata:1968}).

As is known since \cite{E:Vas:nongroup_perfect},
the number of nonequivalent perfect codes grows double-exponentially with respect to the length
(a survey of some constructions can be found in \cite{E:Sol:2008:survey},
\cite{E:Hed:2008:survey}).
Before \cite{E:Vas:nongroup_perfect},
only the linear perfect code, the Hamming code, was known;
moreover because of some strict metrical invariant properties of the perfect codes
\cite{E:Lloyd}, \cite{E:ShSl}, it was conjectured that there are no other
$1$-error correcting perfect binary codes.

At the moment, it is known only two classes of Preparata codes.
The first class contains the original Preparata codes \cite{E:Preparata:1968}
and their generalizations
\cite{E:Dumer:1976}, \cite{E:BvLW}, \cite{E:vDamFDFla:2000}
(the last paper contains the most general known representation
of codes from this class).
These codes are nonlinear, but the including perfect code is linear, the Hamming code.
The second class is the class of $Z_4$-linear Preparata codes.
Strictly speaking, the corresponding
distance-$6$ extended Preparata codes are $Z_4$-linear.
The first known series of such codes was published in \cite{E:HammonsOth:Z4_linearity}.
As was shown in \cite{E:Kantor:2004symplecticsemifield},
the number of nonequivalent $Z_4$-linear extended Preparata codes of length
 $n=4^m$ grows over-polynomially in $n$ for almost all values of $m$.
All Preparata codes from the second class are includes in the same, up to equivalence,
perfect code, whose extension is $Z_4$-linear.
This perfect code is nonlinear if $n>15$ \cite{E:HammonsOth:Z4_linearity}.

So, for every length $n=4^m-1>15$, only two nonequivalent perfect codes
including Preparata codes are known.
The fraction of these codes among the set of all perfect codes is unknown.
It is naturally to conjecture that these fraction
is negligibly small; however,
before the current work,
there was no perfect code known that has length $4^m-1 \ge 63$
and guaranteedly does not include any
Preparata code.
In the current correspondence, we show a short proof of the existence of such perfect codes,
also hoping that this will attract an attention to the problem and stimulate
obtaining more essential results in this direction.

Note that for the length $15$ the situation is clear
(although it cannot be used to predict the asymptotic behavior).
The Preparata code of length $15$ is unique up to equivalence
and known as the Nordstrom-Robinson code, see, e.g., \cite{E:MWS};
this code is included in the Hamming code.
Hence, all nonlinear perfect codes of length $15$ do not include Preparata codes.

\section{Preliminaries}\label{Prelim}

We study codes in the $n$-\emph{dimensional binary Hamming space},
consisting from the set $Q_n$ of all binary $n$-tuples (words), with component-wise
modulo-$2$ addition and the Hamming metric.
The \emph{support} $\mathrm{supp}(\alpha)$ of the word $\alpha$ is the set of its nonzero positions;
the cardinality of the support of a word  $\alpha$ is its \emph{Hamming weight} $\mathrm{wt}(\alpha)$.
The  \emph{Hamming distance} $\rho(\alpha,\beta)$ between words $\alpha$ and $\beta$
is the Hamming weight of $\alpha+\beta$.
By $|\alpha|$, we denote the modulo-$2$ sum of the coordinates of the word $\alpha$.
For binary words $\alpha$ and $\beta$ (their lengths may be different),
$(\alpha,\beta)$ denotes their concatenation.

A set  $C\subseteq Q^n$ of $M$ words with mutual distance at least  $d$
is called a \emph{binary $(n,M,d)$ code}, i.e., a code of length $n$, size $M$, and distance $d$.
A code is called \emph{perfect (with distance $3$)} if the balls of radius $1$ centered
in the code words do not intersect and cover all $Q_n$.
It is straightforward from the definition that the minimal distance between codewords is $3$.
Perfect codes of length $n$ exist for every $n$ of form $n=2^t-1$
and do not exist for any other $n$ (because the cardinality of a radius-$1$ ball must divide the cardinality of $Q_n$).
For every $n=2^t-1$,
there exist a unique,
up to equivalence, linear
(i.e., closed with respect to addition) perfect code, the Hamming code.
In the half of the cases, namely, when $t$ is even,
there are Preparata codes of length $n=2^t-1$,
which are defined as the codes of distance $5$ and size $2^{n+1}/(n+1)^2$.
Every Preparata code is included in a unique perfect code~\cite{E:ZZS:73:Praparata}.

A \emph{Steiner triple system} of order $n$, or $\mathrm{STS}(n)$,
is a collection of $3$-subsets of $\{1,\ldots,n\}$, called \emph{blocks},
such that every pair of elements are included in exactly one block.
The cardinality of $\mathrm{STS}(n)$ is $n(n-1)/6$.
Given a perfect code $C$ and its codeword $\alpha$, 
the set of words $\beta$ of $C$ at distance $3$
from  $\alpha$ defines the Steiner triple system $T(\alpha)$
as follows:
$$T(\alpha) = \{\mathrm{supp}(\beta+\alpha) \ : \ \beta\in C, \ \rho(\alpha,\beta)=3\} .$$
If $C$ includes a Preparata code $P$ and $\alpha\in C\backslash P$, 
then this $\mathrm{STS}(n)$ has the subset
\begin{equation}\label{TPrep}
\{\mathrm{supp}(\beta+\alpha) \ : \ \beta\in P, \ \rho(\alpha,\beta)=3\},
\end{equation}
which partitions the set $\{1,\ldots,n\}$ into $3$-subsets \cite{E:ZZS:73:Praparata,SZZ:1971:UPC}.

A nonempty 
subset $R$ of a code $C$ if called an $i$-\emph{component} of $C$
if the code $(C\setminus R)\cup  (R+ \boldsymbol{e}^i)$,
where $\boldsymbol{e}^i$ is a word of weight $1$ with the $i$th nonzero coordinate, has the same parameters as $C$,
and no proper subset of $R$ satisfies this property.
Two words of a distance-$d$ code are called
$i$-\emph{close} if they differ in $d$ positions including $i$.
If an $i$-component $R$ contains a word $\alpha$, then $R$ contains all words $i$-close to $\alpha$
\cite{E:Sol88en}.
The following fact was proven in \cite{E:Tok:2004.en}.
If a perfect code $C$ includes a Preparata code $P$
and  $R$ is an $i$-component of $C$,
then
the set $P\cap R$ is a perfect code in the graph $(R,E)$,
where $E$ is the set of pairs of words from $R$ at distance $3$.

Let us describe the Vasil'ev construction \cite{E:Vas:nongroup_perfect}
of perfect codes.
Let $H_k$ be a Hamming code of length $k=2^t-1$,
and let $\lambda$ be an arbitrary function from $H_k$ to $\{0,1\}$;
then the set
$$ \{ ( \alpha,\alpha+\beta,|\alpha|+\lambda(\beta) ) \ : \ \alpha\in Q_k, \ \beta\in H_k\} $$
is a perfect code of length $n=2k+1$.
In the case $\lambda\equiv 0$, the resulting code is the Hamming code of length $n=2k+1$.

From this representation, we can see the following.
For every word $(\alpha,\alpha+\beta,|\alpha| )$ from the Hamming code,
the $n$-close words are the words 
$(\alpha,\alpha+\beta,|\alpha|+1)+\boldsymbol{e}^j+\boldsymbol{e}^{k+j}$, $j=1,\ldots,k$.
Hence, for every word $\beta$ from $H_k$, the set
\begin{equation}\label{R-comp}
R=R(\beta) = \{ ( \alpha,\alpha+\beta,|\alpha| ) \ : \ \alpha\in Q_k\}
\end{equation}
is an $n$-component (called \emph{linear}).
The set $\{R(\beta) \,:\, \beta \in H_k\}$ of all such components forms a partition of the Hamming code of length $n$.
Each of these $n$-components can be independently replaced by its switching
$R'= R+\boldsymbol{e}^n$;
as a result we can obtain a large number of new perfect codes.

\section{The main result}\label{MainRes}

We are going to show that some Vasil'ev codes,
namely, the codes obtained from the Hamming code
by switching one component,
do not include any Preparata codes.
\begin{theorem} \label{nPr}
Let $n=4^t-1$ and $\beta\in H_{(n-1)/2}$.
Then, no Preparata codes are included in the perfect code
\begin{equation}\label{C}
C = C(\beta)= (H_n\setminus R)\cup R' ,
\end{equation}
where the component $R=R(\beta)$ is defined in (\ref{R-comp})
and $R'=R+ \boldsymbol{e}^n$ is its switching.
\end{theorem}

{\bf Proof}.
Assume that there exists a Preparata code $P\subseteq C$.
Let a word $\boldsymbol{x}$ of $P$ belongs to the ``switched'' component; i.e.,
$\boldsymbol{x}\in P\cap R'$.
Such a word exists because a Preparata code has a nonempty intersection
with each $i$-component of the including perfect code
(see, for ex., \cite{E:Tok:2004.en}; this also follows from the uniqueness of a perfect code including $P$).
At distance $3$ from $\boldsymbol{x}$, we choose a codeword
$\boldsymbol{y}$ of $C$ that
does not belong to the switched component,
$\boldsymbol{y}\in C\setminus R'$.
Such $\boldsymbol{y}$ exists because there are $n(n-1)/6$
codewords at distance $3$
from $\boldsymbol{x}$, but only $(n-1)/2$ of them
are in the same $n$-component as $\boldsymbol{x}$.
Note that $\boldsymbol{y} \not\in P$.

Without loss of generality we can assume that
$\boldsymbol{y} = \boldsymbol{0}=(0,\ldots,0)$
(if it is not the case, we consider $C+\boldsymbol{y}$ instead of $C$ and $P+\boldsymbol{y}$ instead of $P$).
Then $\mathrm{wt}(\boldsymbol{x})=3$;
let $\{i,j,k\}$ be the support of $\boldsymbol{x}$.
For $R'=R(\beta)+ \boldsymbol{e}^n$ to contain
$\boldsymbol{x}$,
the condition
$\mathrm{wt}(\beta)=3$ should be satisfied (see (\ref{R-comp})).
In this case, it is easy to see from (\ref{R-comp}) that $R'$ contains only $4$ words of weight $3$, with the supports
\begin{equation}\label{STS_C}
\{i,j,k\}, \ \{i\,j'\,k'\}, \ \{i',j,k'\}, \ \{i'\,j'\,k\},
\end{equation}
 for some different
 $i'$, $j'$, $k'$ from $\{1\ldots,n\}\setminus\{i,j,k\}$.
The triple $\{i,j,k\}$
belongs to the partition $T$
of $\{1,\ldots,n\}$
into triples corresponding to the weight-$3$ words of the Preparata code $P$,
see (\ref{TPrep});
the other three triples from
(\ref{STS_C}), obviously, do not belong to $T$.

Let us consider the Steiner triple system $S_{H_n}$
corresponding to the weight-$3$ words of the Hamming code $H_n$.
Since $$H_n=(C\setminus R')\cup R,$$
we see that the Steiner triple system $S_{H_n}$ is obtained from the Steiner triple system $S_C$ of $C$ by replacing the triples (\ref{STS_C}),
corresponding to words of  $R'$, by the triples
$$\{i,j,k'\}, \ \{i,j',k\}, \ \{i',j,k\}, \ \{i',j',k'\},$$
corresponding to words of  $R'$.
So, we have $T\setminus\{\{i,j,k\}\}\subseteq S_{H_n}$,
and the words with the supports
from $T\setminus\{\{i,j,k\}\}$ belong to the Hamming code.
As the Hamming code is linear, we have
$$\boldsymbol{u} = \sum_{\mathrm{supp}(\boldsymbol{z})\in T\setminus \{\{i,j,k\}\}} \boldsymbol{z} \in H_n .$$
Since a perfect code is antipodal and
$\boldsymbol{y} = \boldsymbol{0}=(0,\ldots,0)\in H_n$,
we also have $\boldsymbol{1}=(1,\ldots,1)\in H_n$;
hence,
$\boldsymbol{v}=\boldsymbol{1}+\boldsymbol{u}\in H_n$.
However, from the definition of $\boldsymbol{u}$ we get
$\mathrm{supp}(\boldsymbol{v})= \{i,j,k\}$ and, consequently,
$\boldsymbol{v}=\boldsymbol{x}\notin H_n$.
A contradiction.

\section{Conclusion}

We have shown that some Vasil'ev codes do not include Preparata codes.
Taking into account a local character of the proof,
we can conclude that the same is true for a large number of nonequivalent
Vasil'ev codes.
This way does not yield directly that almost all
 Vasil'ev codes do not include Preparata codes;
 however, it is natural to conjecture this.
 We finish with two conjectures based on common properties of the two perfect codes
 that are known to include Preparata codes.

 \begin{conj}
  All the codewords of weight $3$ ($4$) of a (extended) perfect code including a Preparata code form a linear Steiner triple (respectively, quadruple) system,
  i.e., a system equivalent to that is formed by the weight-$3$ (weight-$3$) words of the linear Hamming code.
 \end{conj}
 \begin{conj}
  All $i$-components (for every $i$) of a perfect code including a Preparata code are linear, i.e., equivalent to the set $\{(x,x,|x|)\}$.
 \end{conj}


\providecommand\href[2]{#2} \providecommand\url[1]{\href{#1}{#1}}
  \def\DOI#1{{\small {DOI}:
  \href{http://dx.doi.org/#1}{#1}}}\def\DOIURL#1#2{{\small{DOI}:
  \href{http://dx.doi.org/#2}{#1}}}

\end{document}